# VisualPredicator: Learning Abstract World Models with Neuro-Symbolic Predicates for Robot Planning

### v1

@@ 3 Neuro-Symbolic Predicates
Neuro-Symbolic Predicates (NSPs) represent visually grounded yet logically rich abstractions that enable efficient planning and problem solving. As Figure 2 illustrates, these predicates are neuro-symbolic because they combine programming language constructs (conditionals, numerics, loops and recursion) with API calls to neural vision-language models for evaluating visually-grounded natural language assertions. NSPs can be grounded in visual perception, and also in proprioceptive and object-tracking features, such as object poses, common in robotics (Kumar et al., 2024; 2023b; Curtis et al., 2022; 2024b). We consider two classes of NSPs: primitive and derived. Primitive NSPs are evaluated directly on the raw state, such as Holding(obj) (which would use VLM queries) or GripperOpen (which would use proprioception). Derived NSPs instead determine their truth value based on the truth value of other NSPs, analogous to derived predicates in planning (Thiébaux et al., 2005; McDermott et al., 1998).
@@
Evaluating Primitive NSPs. No VLM is $100\%$ accurate, even for simple queries like “is the robot holding the jug?”, especially in partially observable environments. To increase the accuracy and precision of NSPs, we take the following two measures.
@@ 5 Learning an Abstract World Model from Interacting with the Environment
3:for i $\in$ range($1,n_{\text{max\_ite}}$do
4:     $\mathcal{D}_{i},\rho_{i},\nu_{i}\leftarrow$ Explore($\Psi_{i-1},\Omega_{i-1},\mathcal{E},\mathcal{T}$$\triangleright$ section 5.1
5:     if $\rho_{i}>\rho_{\text{best}}$ or ($\rho_{i}=\rho_{\text{best}}$ and $\nu_{i}<\nu_{\text{best}}$then
6:         $\Psi_{\text{best}},\Omega_{\text{best}},\rho_{\text{best}},\nu_{\text{best}}\leftarrow\Psi_{i},\Omega_{i},\rho_{i},\nu_{i}$
7:     if $\nu_{i}=0$ then
8:         break
9:     $\mathcal{D}\leftarrow\mathcal{D}\cup\mathcal{D}_{i}$
10:     $\Psi^{\prime}\leftarrow\emptyset$
@@
12:         $\Psi^{\prime}\leftarrow$ Propose Predicates($\mathcal{D},\Psi_{i-1}$) $\triangleright$ section 5.2
@@ 5.2 Proposing Predicates
We introduce three strategies for prompting VLMs to invent predicates – two that are conditioned on collected data, and one that is not (see section B.3 for further details).
@@ 5.3 Selecting a Predicate Set
To address this, we devise a novel objective that scores a set of predicates $\Psi$ based on classification accuracy, plus a simplicity bias. The classification score is obtained by first learning HLAs using the set of predicates $\Psi$ (discussed more in section 5.4), and then computing the classification accuracy of the HLAs (see Section B.2). Later in learning, after discovering enough (a hyperparameter one can choose) satisficing plans, we switch to the objective of Silver et al. (2023), which takes planning efficiency and simplicity into account.
We perform a greedy best-first search with either score function as the heuristic. It starts from the set of goal predicates $\Psi_{G}$ and adds a single new predicate from the proposed candidates at each step, and finally returns the set of predicates with the highest score.
@@ 6 Experiments
Experimental Setup. We evaluated seven different approaches across five robotic environments simulated using the PyBullet physics engine (Coumans & Bai, 2016). Each result is averaged over five random seeds, and for each seed, we sample 50 test tasks that feature more objects and more complex goals than those encountered during training. The agent is provided with 5 training tasks in the Cover and Coffee environments, 10 tasks in Cover Heavy and Balance, and 20 tasks in Blocks. The planning budget $n_{\text{abstract}}$ is set to 8 for all domains except Coffee, where it is set to 100.
Environments. We briefly discuss the environments used, with more details in appendix D.
@@ 8 Conclusion
In this work, we introduced Neuro-Symbolic Predicates (NSPs), a novel representation that combines the flexibility of neural networks to represent open-ended, visually grounded concepts, and the interpretability and compositionality of symbolic representations, for planning. To support this, we developed an online algorithm for inventing NSPs and learning abstract world models, which allows efficient acquisition of NSPs. Our experiments across five simulated robotic domains demonstrated that our method outperforms existing approaches, including hierarchical reinforcement learning, VLM planning, and traditional symbolic predicates, particularly in terms of sample efficiency, generalization, and interpretability. Future work will focus on incorporating recovery mechanisms for failed plans, relaxing assumptions about skills, enhancing exploration efficiency, and scaling to partially observable and real-world domains.
@@ Appendix B Additional Details about the Online Invention Algorithm
B.1 Learning HLAs by extending the cluster-and-intersect algorithm
@@
B.2 Classification-Accuracy-Based Predicate Sets Score Function
@@ B.3 Prompting for predicates
B.4 Limitations and Failure Cases
A primary limitation of the system is the accuracy and reliability of the VLM in evaluating NSPs. For instance, in the Cover Heavy domain, our initial plan was to assign common materials, such as wood and metal, to blocks to distinguish between light and heavy objects. While the predicate proposal VLM successfully suggested appropriate predicates (e.g., IsWood(?block) and IsMetal(?block)), the predicate evaluation VLM was unable to interpret these predicates with sufficient accuracy and consistency to build a useful world model. The issue persisted even after switching to white and black blocks to represent heavy and light blocks and was ultimately resolved by using green and red blocks instead. Similarly, in the Coffee domain, the predicate IsJugFilled(?jug) is an essential precondition for the pour HLA. However, the VLM could not interpret this predicate accurately enough, necessitating that we treat it as a predefined predicate. One potential solution could be to integrate proprioception more effectively into the system.
@@ C.2.1 Cover
NSRT-Op0:
Parameters: [?x0:block, ?x1:robot]
Preconditions: [GripperOpen(?x1:robot)]
Add Effects: [Holding(?x1:robot, ?x0:block)]
Delete Effects: [GripperOpen(?x1:robot)]
Ignore Effects: []
Option Spec: Pick(?x0:block)
NSRT-Op1:
Parameters: [?x0:block, ?x1:robot, ?x2:target]
Preconditions: [Holding(?x1:robot, ?x0:block)]
Add Effects: [Covers(?x0:block, ?x2:target), GripperOpen(?x1:robot)]
Delete Effects: [Holding(?x1:robot, ?x0:block)]
Ignore Effects: []
Option Spec: Place(?x0:block, ?x2:target)
@@ C.2.2 Blocks
NSRT-Op0:
Parameters: [?x0:block, ?x1:block, ?x2:robot]
Preconditions: [Clear(?x1:block), EmptyGripper(?x2:robot), On(?x1:block, ?x0:block)]
Add Effects: [Gripping(?x2:robot, ?x1:block)]
Delete Effects: [EmptyGripper(?x2:robot), On(?x1:block, ?x0:block)]
Ignore Effects: []
Option Spec: Pick(?x2:robot, ?x1:block)
NSRT-Op1:
Parameters: [?x0:block, ?x1:robot]
Preconditions: [Gripping(?x1:robot, ?x0:block)]
Add Effects: [EmptyGripper(?x1:robot), OnTable(?x0:block)]
Delete Effects: [Gripping(?x1:robot, ?x0:block)]
Ignore Effects: []
Option Spec: PutOnTable(?x1:robot)
NSRT-Op2:
Parameters: [?x0:block, ?x1:robot]
Preconditions: [Clear(?x0:block), EmptyGripper(?x1:robot), OnTable(?x0:block)]
Add Effects: [Gripping(?x1:robot, ?x0:block)]
Delete Effects: [EmptyGripper(?x1:robot), OnTable(?x0:block)]
Ignore Effects: []
Option Spec: Pick(?x1:robot, ?x0:block)
NSRT-Op3:
Parameters: [?x0:block, ?x1:block, ?x2:robot]
Preconditions: [Clear(?x0:block), Gripping(?x2:robot, ?x1:block)]
Add Effects: [EmptyGripper(?x2:robot), On(?x1:block, ?x0:block)]
Delete Effects: [Gripping(?x2:robot, ?x1:block)]
Ignore Effects: []
Option Spec: Stack(?x2:robot, ?x0:block)
@@ C.2.3 Coffee
NSRT-Op0:
Parameters: [?x0:jug, ?x1:robot]
Preconditions: [GripperOpen(?x1:robot), JugUpright(?x0:jug)]
Add Effects: [RobotHoldingJug(?x1:robot, ?x0:jug)]
Delete Effects: [GripperOpen(?x1:robot)]
Ignore Effects: []
Option Spec: PickJug(?x1:robot, ?x0:jug)
NSRT-Op1:
Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot]
Preconditions: [RobotHoldingJug(?x2:robot, ?x1:jug)]
Add Effects: [GripperOpen(?x2:robot), JugInMachine(?x1:jug, ?x0:coffee_machine)]
Delete Effects: [RobotHoldingJug(?x2:robot, ?x1:jug)]
Ignore Effects: []
Option Spec: PlaceJugInMachine(?x2:robot, ?x1:jug, ?x0:coffee_machine)
NSRT-Op2:
Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot]
Preconditions: [JugInMachine(?x1:jug, ?x0:coffee_machine)]
Add Effects: [JugFilled(?x1:jug)]
Delete Effects: []
Ignore Effects: []
Option Spec: TurnMachineOn(?x2:robot, ?x0:coffee_machine)
NSRT-Op3:
Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot]
Preconditions: [JugInMachine(?x1:jug, ?x0:coffee_machine)]
Add Effects: [RobotHoldingJug(?x2:robot, ?x1:jug)]
Delete Effects: [GripperOpen(?x2:robot), JugInMachine(?x1:jug, ?x0:coffee_machine)]
Ignore Effects: []
Option Spec: PickJug(?x2:robot, ?x1:jug)
NSRT-Op4:
Parameters: [?x0:cup, ?x1:jug, ?x2:robot]
Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)]
Add Effects: [CupFilled(?x0:cup)]
Delete Effects: [JugFilled(?x1:jug), JugUpright(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)]
Ignore Effects: []
Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup)
NSRT-Op5:
Parameters: [?x0:jug, ?x1:robot]
Preconditions: [GripperOpen(?x1:robot)]
Add Effects: [JugUpright(?x0:jug)]
Delete Effects: []
Ignore Effects: []
Option Spec: Twist(?x1:robot, ?x0:jug)
NSRT-Op6:
Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot]
Preconditions: [JugInMachine(?x1:jug, ?x0:coffee_machine)]
Add Effects: [JugFilled(?x1:jug)]
Delete Effects: [JugInMachine(?x1:jug, ?x0:coffee_machine)]
Ignore Effects: []
Option Spec: TurnMachineOn(?x2:robot, ?x0:coffee_machine)
NSRT-Op7:
Parameters: [?x0:cup, ?x1:jug, ?x2:robot]
Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)]
Add Effects: [CupFilled(?x0:cup), JugTilted(?x1:jug)]
Delete Effects: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)]
Ignore Effects: []
Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup)
NSRT-Op8:
Parameters: [?x0:cup, ?x1:jug, ?x2:robot]
Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)]
Add Effects: [CupFilled(?x0:cup), JugTilted(?x1:jug)]
Delete Effects: []
Ignore Effects: []
Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup)
NSRT-Op9:
Parameters: [?x0:cup, ?x1:jug, ?x2:robot]
Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)]
Add Effects: [CupFilled(?x0:cup), JugTilted(?x1:jug)]
Delete Effects: [RobotHoldingJug(?x2:robot, ?x1:jug)]
Ignore Effects: []
Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup)

### v2

@@ 3 Neuro-Symbolic Predicates
Neuro-Symbolic Predicates (NSPs) represent visually grounded yet logically rich abstractions that enable efficient planning and problem solving. As Figure 2 illustrates, these predicates are neuro-symbolic because they combine programming language constructs (conditionals, numerics, loops and recursion) with API calls to neural vision-language models for evaluating visually-grounded natural language assertions. NSPs can be grounded in visual perception, and also in proprioceptive and object-tracking features, such as object poses, common in robotics (Kumar et al., 2024; 2023b; Curtis et al., 2022; 2024b). We consider two classes of NSPs: primitive and derived. Primitive NSPs are evaluated directly on the raw state, such as Holding(obj) (which would use VLM queries) or GripperOpen() (which would use proprioception). Derived NSPs instead determine their truth value based on the truth value of other NSPs, analogous to derived predicates in planning (Thiébaux et al., 2005; McDermott et al., 1998).
@@
Evaluating Primitive NSPs. No VLM is $100\%$ accurate, even for simple queries like “is the robot holding the jug?”, especially in partially observable environments. To increase the accuracy and precision of NSPs, we take the following two measures (see Section B.1 for an example prompt).
@@ 5 Learning an Abstract World Model from Interacting with the Environment
3:init: $\Psi^{\prime}\leftarrow\Psi_{0}$
4:for i $\in$ range($1,n_{\text{max\_ite}}$do
5:     $\mathcal{D}_{i},\rho_{i},\nu_{i}\leftarrow$ Explore($\Psi_{i-1},\Omega_{i-1},\mathcal{E},\mathcal{T}$$\triangleright$ section 5.1
6:     if $\rho_{i}>\rho_{\text{best}}$ or ($\rho_{i}=\rho_{\text{best}}$ and $\nu_{i}<\nu_{\text{best}}$then
7:         $\Psi_{\text{best}},\Omega_{\text{best}},\rho_{\text{best}},\nu_{\text{best}}\leftarrow\Psi_{i},\Omega_{i},\rho_{i},\nu_{i}$
8:     if $\nu_{i}=0$ then
9:         break
10:     $\mathcal{D}\leftarrow\mathcal{D}\cup\mathcal{D}_{i}$
@@
12:         $\Psi^{\prime}\leftarrow\Psi^{\prime}\cup$ Propose NSPs($\mathcal{D},\Psi_{i-1}$) $\triangleright$ section 5.2
@@ 5.2 Proposing Predicates
We introduce three strategies for prompting VLMs to invent diverse, task-relevant predicates – two that are conditioned on collected data, and one that is not (see Section B.4 for further details).
@@ 5.3 Selecting a Predicate Set
To address this, we devise a novel objective that scores a set of predicates $\Psi$ based on classification accuracy, plus a simplicity bias. The objective score is obtained by first learning HLAs using the set of predicates $\Psi$ (discussed more in section 5.4), and then computing the classification accuracy of the HLAs (see Section B.3). Later in learning, after discovering enough (a hyperparameter one can choose) satisficing plans, we switch to the objective of Silver et al. (2023), which takes planning efficiency and simplicity into account.
We perform a greedy best-first search (GBFS) with either score function as the heuristic. It starts from the set of goal predicates $\Psi_{G}$ and adds a single new predicate from the proposed candidates at each step, and finally returns the set of predicates with the highest score. This effectively selects $\Psi_{i}\leftarrow\arg\min_{\Psi\subseteq\mathcal{P}(\Psi)}{J}(\Psi^{\prime})$, where $J(\cdot)$ is the objective function. In our experiments, we found the search space is small enough that the enumeration typically takes just a few minutes on a single CPU. For larger search spaces, local hill climbing could be used in place of GBFS.
@@ 6 Experiments
Experimental Setup. We evaluated seven different approaches across five robotic environments simulated using the PyBullet physics engine (Coumans & Bai, 2016). Each result is averaged over five random seeds, and for each seed, we sample 50 test tasks that feature more objects and more complex goals than those encountered during training. The agent is provided with 5 training tasks in the Cover and Coffee environments, 10 tasks in Cover Heavy and Balance, and 20 tasks in Blocks. The planning budget $n_{\text{abstract}}$ is set to 8 for all domains except Coffee, where it is set to 100. The approaches are run on a single CPU except for the MAPLE baseline, which utilizes uses a GPU.
Environments. We briefly describe the environments used, including their hand-coded closed-loop controllers, which are shared across all approaches. Additional details can be found in appendix D.
@@ 8 Conclusion
In this work, we introduced Neuro-Symbolic Predicates (NSPs), a novel representation that combines the flexibility of neural networks to represent open-ended, visually grounded concepts, and the interpretability and compositionality of symbolic representations, for planning. To support this, we developed an online algorithm for inventing NSPs and learning abstract world models, which allows efficient acquisition of NSPs. Our experiments across five simulated robotic domains demonstrated that our method outperforms existing approaches, including hierarchical reinforcement learning, VLM planning, and traditional symbolic predicates, particularly in terms of sample efficiency, generalization, and interpretability. Exciting areas for future work include incorporating recovery mechanisms for failed plans, enhancing exploration efficiency, scaling to partially observable and real-world domains, and relaxing assumptions about skills leverage advances in policy synthesis (Liang et al., 2023), RL (Liang et al., 2024; Ma et al., 2023), or motion planning (Huang et al., 2024).
@@ Appendix B Additional Details about the Online Invention Algorithm
B.1 Predicate Interpretation
We provide an example prompt used to interpret the truth value of the ground atom DirectlyOn(block5, block6) in the state with cropped observation shown in Figure 5. The highlighted text illustrates how we condition on previous action, previous observation, previous truth value, and object IDs to improve the predicate evaluation accuracy.
Evaluate the truth value of the following assertions in the current state as depicted by the image labeled with ’curr. state’. For context, the state is right after the robot has successfully executed the action Pick(robot1:robot, block5:block). The state before executing the action is depicted by the image labeled with ’prev. state’. Please carefully examine the images depicting the ’prev. state’ and ’curr. state’ before making a judgment. The assertions to evaluate are: 1. block5 is directly on top of block6. (which was False before the successful execution of the previous action)
Refer to caption Refer to caption
Figure 5: Example cropped current (right) and previous (left) observations used for interpreting ground predicates.
B.2 Learning HLAs by extending the cluster and intersect algorithm
@@
In terms of time complexity, cluster and intersect is linear in the number of successful transitions in $\mathcal{D}$ and the number of predicates $\Psi$, where the additional greedy best-first search (GBFS) that we do introduces exponential complexity with respect to the number of predicates. To balance computational efficiency and performance, we use cluster and intersect in the inner loop of predicate section and then apply our method to the selected predicates (which is usually less than a dozen). Additionally, local hill climbing can be used as an alternative to GBFS to further improve computational efficiency.
B.3 Classification-Accuracy-Based Predicate Sets Score Function
@@ B.4 Prompting for predicates
B.5 Limitations and Failure Cases
A primary limitation of the system is the accuracy and reliability of the VLM in evaluating NSPs.
In some cases, the system can recover from imperfect predicate evaluation accuracies. This is because noisy predicates are not selected during the predicate selection process, and variations of the predicates, with slightly different natural language descriptions, can be proposed in later invention iterations. These variations may achieve higher scores, making them more likely to be selected.
In other cases, they never recover. For instance, in the Cover Heavy domain, our initial plan was to assign common materials, such as wood and metal, to blocks to distinguish between light and heavy objects. While the predicate proposal VLM successfully suggested appropriate predicates (e.g., IsWood(?block) and IsMetal(?block)), the predicate evaluation VLM was unable to interpret these predicates with sufficient accuracy and consistency to build a useful world model. The issue persisted even after switching to white and black blocks to represent heavy and light blocks and was ultimately resolved by using green and red blocks instead. Similarly, in the Coffee domain, the predicate IsJugFilled(?jug) is an essential precondition for the pour HLA. However, the VLM could not interpret this predicate accurately enough, necessitating that we treat it as a predefined predicate.
Potential solutions include: 1) integrating proprioception more effectively into the system; 2) developing ways to accurately assign belief scores over the truth values (e.g., use “IsJugFilled(?jug)–0.9” to denote “I believe the jug is filled with coffee with probability 0.9”); Or designing embeddings for ground predicates and observations, and determining the truth values of ground predicates by comparing distances between the corresponding embeddings.
At the same time, we expect improved accuracy in real-world scenarios compared to simulated domains with poor graphics quality, as there should be less distribution shift relative to the VLMs’ training data, and the VLMs have demonstrated very strong performance on simple visual question-answering tasks with natural images Yang et al. (2023).
@@ C.2.1 Cover
NSRT-Op0: Parameters: [?x0:block, ?x1:robot] Preconditions: [GripperOpen(?x1:robot)] Add Effects: [Holding(?x1:robot, ?x0:block)] Delete Effects: [GripperOpen(?x1:robot)] Ignore Effects: [] Option Spec: Pick(?x0:block) NSRT-Op1: Parameters: [?x0:block, ?x1:robot, ?x2:target] Preconditions: [Holding(?x1:robot, ?x0:block)] Add Effects: [Covers(?x0:block, ?x2:target), GripperOpen(?x1:robot)] Delete Effects: [Holding(?x1:robot, ?x0:block)] Ignore Effects: [] Option Spec: Place(?x0:block, ?x2:target)
@@ C.2.2 Blocks
NSRT-Op0: Parameters: [?x0:block, ?x1:block, ?x2:robot] Preconditions: [Clear(?x1:block), EmptyGripper(?x2:robot), On(?x1:block, ?x0:block)] Add Effects: [Gripping(?x2:robot, ?x1:block)] Delete Effects: [EmptyGripper(?x2:robot), On(?x1:block, ?x0:block)] Ignore Effects: [] Option Spec: Pick(?x2:robot, ?x1:block) NSRT-Op1: Parameters: [?x0:block, ?x1:robot] Preconditions: [Gripping(?x1:robot, ?x0:block)] Add Effects: [EmptyGripper(?x1:robot), OnTable(?x0:block)] Delete Effects: [Gripping(?x1:robot, ?x0:block)] Ignore Effects: [] Option Spec: PutOnTable(?x1:robot) NSRT-Op2: Parameters: [?x0:block, ?x1:robot] Preconditions: [Clear(?x0:block), EmptyGripper(?x1:robot), OnTable(?x0:block)] Add Effects: [Gripping(?x1:robot, ?x0:block)] Delete Effects: [EmptyGripper(?x1:robot), OnTable(?x0:block)] Ignore Effects: [] Option Spec: Pick(?x1:robot, ?x0:block) NSRT-Op3: Parameters: [?x0:block, ?x1:block, ?x2:robot] Preconditions: [Clear(?x0:block), Gripping(?x2:robot, ?x1:block)] Add Effects: [EmptyGripper(?x2:robot), On(?x1:block, ?x0:block)] Delete Effects: [Gripping(?x2:robot, ?x1:block)] Ignore Effects: [] Option Spec: Stack(?x2:robot, ?x0:block)
@@ C.2.3 Coffee
NSRT-Op0: Parameters: [?x0:jug, ?x1:robot] Preconditions: [GripperOpen(?x1:robot), JugUpright(?x0:jug)] Add Effects: [RobotHoldingJug(?x1:robot, ?x0:jug)] Delete Effects: [GripperOpen(?x1:robot)] Ignore Effects: [] Option Spec: PickJug(?x1:robot, ?x0:jug) NSRT-Op1: Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot] Preconditions: [RobotHoldingJug(?x2:robot, ?x1:jug)] Add Effects: [GripperOpen(?x2:robot), JugInMachine(?x1:jug, ?x0:coffee_machine)] Delete Effects: [RobotHoldingJug(?x2:robot, ?x1:jug)] Ignore Effects: [] Option Spec: PlaceJugInMachine(?x2:robot, ?x1:jug, ?x0:coffee_machine) NSRT-Op2: Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot] Preconditions: [JugInMachine(?x1:jug, ?x0:coffee_machine)] Add Effects: [JugFilled(?x1:jug)] Delete Effects: [] Ignore Effects: [] Option Spec: TurnMachineOn(?x2:robot, ?x0:coffee_machine) NSRT-Op3: Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot] Preconditions: [JugInMachine(?x1:jug, ?x0:coffee_machine)] Add Effects: [RobotHoldingJug(?x2:robot, ?x1:jug)] Delete Effects: [GripperOpen(?x2:robot), JugInMachine(?x1:jug, ?x0:coffee_machine)] Ignore Effects: [] Option Spec: PickJug(?x2:robot, ?x1:jug) NSRT-Op4: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)] Add Effects: [CupFilled(?x0:cup)] Delete Effects: [JugFilled(?x1:jug), JugUpright(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)] Ignore Effects: [] Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup) NSRT-Op5: Parameters: [?x0:jug, ?x1:robot] Preconditions: [GripperOpen(?x1:robot)] Add Effects: [JugUpright(?x0:jug)] Delete Effects: [] Ignore Effects: [] Option Spec: Twist(?x1:robot, ?x0:jug)
NSRT-Op6: Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot] Preconditions: [JugInMachine(?x1:jug, ?x0:coffee_machine)] Add Effects: [JugFilled(?x1:jug)] Delete Effects: [JugInMachine(?x1:jug, ?x0:coffee_machine)] Ignore Effects: [] Option Spec: TurnMachineOn(?x2:robot, ?x0:coffee_machine) NSRT-Op7: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)] Add Effects: [CupFilled(?x0:cup), JugTilted(?x1:jug)] Delete Effects: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)] Ignore Effects: [] Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup) NSRT-Op8: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)] Add Effects: [CupFilled(?x0:cup), JugTilted(?x1:jug)] Delete Effects: [] Ignore Effects: [] Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup) NSRT-Op9: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [JugFilled(?x1:jug), RobotHoldingJug(?x2:robot, ?x1:jug)] Add Effects: [CupFilled(?x0:cup), JugTilted(?x1:jug)] Delete Effects: [RobotHoldingJug(?x2:robot, ?x1:jug)] Ignore Effects: [] Option Spec: Pour(?x2:robot, ?x1:jug, ?x0:cup)